OpenStudy (anonymous):

what is 8+the square root of 13?

OpenStudy (anonymous):

I don't understand the question. That expression is already in simplest form, what else do you want to do with it?

OpenStudy (anonymous):

use your calculator =o

OpenStudy (anonymous):

i mean you can technically estimate it by finding the closest roots

OpenStudy (anonymous):

@Outkast3r09 , yes, to get a decimal approximation, a calculator is the best way. There are old iterative methods for finding square roots by hand, but why bother? ;-)

OpenStudy (anonymous):

square roots i mean so \[\sqrt 9 < \sqrt 13 < \sqrt 16\] \[3<\sqrt 13 <4\]

OpenStudy (anonymous):

some teachers don't allow calculators... imagine that lol

OpenStudy (anonymous):

Yeah, that's a good way to get a quick mental estimate. I doubt that's what @ihatealgebrasomuch is asking for though.

OpenStudy (anonymous):

;-) I happen to be one of those teachers who doesn't allow calculators, that's why I said the expression is already in simplest form.